OpenStudy (anonymous):

An aircraft, used for fire spotting, flies from its base to locate a fire at an unknown distance, x km away. It travels straight to the fire and back, averaging 240 km/h for the outward trip and 320 km/h for the return trip. If the plane was away for 35 minutes, find the distance, x km. Is there supposed to be some sort of simultaneous equation involved, with the use of speed = d/t or am I supposed to sub all of the information into the one equation. Also, not sure if I need to convert 35mins into hours or not?

OpenStudy (astrophysics):

Yes, convert all the units so they are consistent, and always start by making a diagram.  you have the right idea, we are using v = d/t for the round trip. \(v_1 = 240 km/h\) and \(v_2 = 320 km/h\) and we know that the distance \(x_1~~ \text{and}~~ x_2\) are the same so we have \(x_1 = x_2 = x\) and the time \(t_1+t_2 = 35 \min\) since we know the plane was away for 35 minutes. I think this should be enough information to get your started and figure it out.